OpenStudy (atomicreaper):

Angle θ lies in the fourth quadrant, and cot θ = -8/15. Sin θ= Cos θ=

OpenStudy (er.mohd.amir):

cse^2 x=1+cot^2x =1+(64/225) cse x=-17/15 sin x=-15/17 in IV QRNT and cos^2 x=1-sin^2 x=1-(225/289) cos x=8/17

OpenStudy (atomicreaper):

Thanks